when you give me the answer could you explain it please so that i understand.
It takes 1 hour longer to fly to St. Paul at 200mi/h than it does to return at 250 mi/h. How far away is St. Paul?

Answer by vasu2qute(17)

let us assume that it takes xhrs to return at 250mi/hr
then time taken to fly to St.Paul = (x+1)hrs at 200m/hr
distance=time x speed
onward distance= 250x
return distance=200(x+1)
onward distance=return distance
250x=200(x+1)
250x=200x+200
250-200x=200
50x=200
x=200/50
x=4hrs
distance=250x=250x4=1000miles